X + Y ATE 52 COOKIES
IF X ATE 11 COOKIES, HE WOULD HAVE EATEN 19 COOKIES LESS THAN Y
HOW MANY COOKIES DID Y HAVE IN THE BEGINNING?

Answer by Maths68(1474)

X + Y ATE 52 COOKIES
IF X ATE 11 COOKIES, HE WOULD HAVE EATEN 19 COOKIES LESS THAN Y
HOW MANY COOKIES DID Y HAVE IN THE BEGINNING?
X ATE = 11
Y ATE = 11+19 = 30
COOKIES LEFT WITH Y =52 - (X ATE + Y ATE) = 52-(11+30) = 52-41=11
cOOKIES THAT Y HAS IN THE BEGINING = Y ATE + COOKIES LEFT = 30+11=41